OpenStudy (anonymous):

The equation of a certain circle is \[x^{2}+y^{2}+6x+6y=87\], find the equation of a circle tangent to the given circle and center at (2,-2).

OpenStudy (anonymous):

Center at (2, -2) Center of given circle is (-2 , -2) Clearly, The distance between the centres is 4 units. Also radii of given circle = \[\sqrt{3^2 + 3^2 + 87} = \sqrt{105}\]

OpenStudy (anonymous):

** Center of given circle is (-3 , -3) And distance of the centers is \[\sqrt{5^2 + 1^2} = \sqrt{26}\]

OpenStudy (anonymous):

Therefore the situation is like :